# Counterintuitive Potential-Barrier Affinity Effect

### v1

@@ Results
We subsequently calculate the effective local potential by the Kohn-Sham density functional theory[20] in Figure 2a, along with the valence band maximum (VBM) and conduction band minimum (CBM). Contrary to the conventional expectation of potential wells, we found that electrons localize at the maximum point of the potential ($V_{M}$) within the interstitial region. Furthermore, we observed that the Fermi energy level ($E_{F}=16.0$ eV) is much higher than the maximum point of the potential ($V_{M}=4.4$ eV), indicating that near-free electrons (NFEs) occupy the single-particle states with energy levels ranging from $V_{M}$ to $E_{F}$. The band structure shown in Figure 2b (left) displays two valence electron bands (4 electrons per cell) located within the NFE energy range. This suggests that all valence electrons behave as unbound-state NFEs rather than being confined as interstitial bound electrons. Additionally, the density of states (DOS) in Figure 2b (right) demonstrates that the s, p, and d components are negligible compared to the total DOS within the valence-electron energy range, which is also a distinct feature commonly associated with unbound-state NFEs[21], rather than the multicentered bonds of hybrid orbitals. To further clearly demonstrate the presence of the electron localization within the interstitial region, we plot the Kohn-Sham effective local potential ($V_{eff}^{KS}$), total ($\rho$), bound-state ($\rho_{B}$), and near-free ($\rho_{F}$) electron densities along the $\bar{1}1$] direction of Na-hP4 in Figure 2c. We observed that the NFE bands contribute almost all the interstitial electron density near the maximum potential point, while the electron density tails from the bound-state bands can be ignored. This result directly confirms that interstitial electrons originate from near-free-valence states. Notably, we further found evidence for the existence of NFEs ($E_{F}>V_{M}$) in representative electrides[22, 23, 8, 24, 25, 26, 27], as illustrated in Figure 2d, suggesting that the existence of near-free valence electrons is a common phenomenon among electrides. These findings compel a reevaluation of the current understanding of IAEs in electrides.
@@
where $\psi^{\text{A}}_{nk}$ and $\psi^{\text{B}}_{nk}$ are the branches of the wavefunction in regions A and B, respectively; $A_{nk}$, $A_{nk}^{\prime}$, $B_{nk}$, and $B_{nk}^{\prime}$ are the corresponding plane-wave-component coefficients in the wavefunction. The wavevectors ($\alpha_{nk}$ and $\beta_{nk}$) in A and B regions are $\alpha_{nk}=\sqrt{2\epsilon_{nk}}$ and $\beta_{nk}=\sqrt{2(\epsilon_{nk}-V_{0})}$, respectively. The details of the numerical solution for the Schrödinger equation of KP model can be found in the Method Section. In Figure 3b, we present the unbound-state band structure ($\epsilon_{nk}>V_{0}$) under the KP potential, which shows finite band gaps and parabolic characteristics as the perturbed free-electron gas[29, 30]. The concavities of the unbound-state bands in Figure 3b are also similar to that in Figure 2b for Na-hP4 near the $G$ point.
Subsequently, we calculated the absolute value of the wavefunctions and the electron density in Figure 3c and Figure 3d, respectively, corresponding to the $n=1,2$ bands in Figure 3b. All of them exhibit counterintuitive phenomena, where the large amplitudes are observed in the interatomic region. This is consistent with the electron density of Na-hP4 in Figure 2c. Furthermore, we reproduced the double peaks of the interatomic electron density in Figure 2c by our unbound-state KP model. Figure 3c shows that the band index ($n$) is directly related to the number of peaks of $|\psi_{nk}(x)|$ in the barrier region. Herein, we do not consider the electron density in the atomic region because of the large differences in the atomic central (pseudo) potential and the square potential. Instead, we will mainly use the unbound-state KP model to demonstrate that the steep-shaped potential can lead to a high electron density in the interatomic region. In the unbound-state regime ($\epsilon_{nk}>V_{0}$), we can directly find that the wavefunction exhibits the properties of short and long wavelengths in regions A (atomic) and B (interatomic) regions, respectively, because of the inequality $\alpha_{nk}>\beta_{nk}$. Therefore, the wavefunction-continuity conditions at the point of barrier-well junction, such as $x=0$, can be written as:
@@
According to the results of the KP model, we find that the PBA of quantum state is dramatically consistent with the classical physical intuition: a free particle has an initial kinetic energy $T_{0}$; When it passes through the barrier region $V(x)>0$, part of the kinetic energy is converted into potential energy, and its velocity decreases due to the smaller kinetic energy $T(x)=T_{0}-V(x)<T_{0}$, causing it to have a higher probability of being detected in the higher barrier region. When the initial kinetic energy is much higher than the potential barrier $T_{0}>>V(x)$, the particle can move freely throughout space $T(x)\approx T_{0}$.

### v2

@@ Results
We subsequently calculate the effective local potential by the Kohn-Sham density functional theory[20] in Figure 2a, along with the valence band maximum (VBM) and conduction band minimum (CBM). Contrary to the conventional expectation of potential wells, we found that electrons localize at the maximum point of the potential ($V_{M}$) within the interstitial region. Furthermore, we observed that the Fermi energy level ($E_{F}=16.0$ eV) is much higher than the maximum point of the potential ($V_{M}=4.4$ eV), indicating that near-free electrons (NFEs) occupy the single-particle states with energy levels ranging from $V_{M}$ to $E_{F}$. The band structure shown in Figure 2b (left) displays two valence electron bands (4 electrons per cell) located within the NFE energy range. This suggests that all valence electrons behave as unbound-state NFEs rather than being confined as interstitial bound electrons. Additionally, the density of states (DOS) in Figure 2b (right) demonstrates that the s, p, and d components are negligible compared to the total DOS within the valence-electron energy range, which is also a distinct feature commonly associated with unbound-state NFEs[21], rather than the multicentered bonding of hybrid orbitals. To further clearly demonstrate the presence of the electron localization within the interstitial region, we plot the Kohn-Sham effective local potential ($V_{eff}^{KS}$), total ($\rho$), bound-state ($\rho_{B}$), and near-free ($\rho_{F}$) electron densities along the $\bar{1}1$] direction of Na-hP4 in Figure 2c. We observed that the NFE bands contribute almost all the interstitial electron density near the maximum potential point, while the electron density tails from the bound-state bands can be ignored. This result directly confirms that interstitial electrons originate from near-free-valence states. We also present the effective ionic potential of Na+-hP4 in Figure 5 of the Method Section by removing the valence electrons. The analysis further confirms that IAEs behave as NFEs with high density accumulation around the potential maximum. Notably, we further found evidence for the existence of NFEs ($E_{F}>V_{M}$) in representative electrides[22, 23, 8, 24, 25, 26, 27], as illustrated in Figure 2d, suggesting that the existence of near-free valence electrons is a common phenomenon among electrides. These findings compel a reevaluation of the current understanding of IAEs in electrides.
@@
where $\psi^{\text{A}}_{nk}$ and $\psi^{\text{B}}_{nk}$ are the branches of the wavefunction in regions A and B, respectively; $A_{nk}$, $A_{nk}^{\prime}$, $B_{nk}$, and $B_{nk}^{\prime}$ are the corresponding plane-wave-component coefficients in the wavefunction. The wavevectors ($\alpha_{nk}$ and $\beta_{nk}$) in A and B regions are $\alpha_{nk}=\sqrt{2\epsilon_{nk}}$ and $\beta_{nk}=\sqrt{2(\epsilon_{nk}-V_{0})}$, respectively. The details of the numerical solution for the Schrödinger equation of KP model can be found in the Method Section. In Figure 3b, we present the unbound-state band structure ($\epsilon_{nk}>V_{0}$) under the KP potential, which shows finite band gaps and parabolic characteristics as the perturbed free-electron gas[29, 30]. The concavities of the unbound-state bands in Figure 3b are also similar to those in Figure 2b for Na-hP4 near the $G$ point.
Subsequently, we calculated the absolute value of the wavefunctions and the electron density in Figure 3c and Figure 3d, respectively, corresponding to the $n=1,2$ bands in Figure 3b. All of them exhibit counterintuitive phenomena, where the large amplitudes are observed in the interatomic region. This is consistent with the electron density of Na-hP4 in Figure 2c. Furthermore, we reproduced the double peaks of the interatomic electron density in Figure 2c by our unbound-state KP model. Figure 3c shows that the band index ($n$) is directly related to the number of peaks of $|\psi_{nk}(x)|$ in the barrier region. Herein, we do not consider the electron density in the atomic region because of the large differences in the atomic central potential and the square potential. Instead, we will mainly use the unbound-state KP model to demonstrate that the steep-shaped potential can lead to a high electron density in the interatomic region. In the unbound-state regime ($\epsilon_{nk}>V_{0}$), we can directly find that the wavefunction exhibits the properties of short and long wavelengths in regions A (atomic) and B (interatomic) regions, respectively, because of the inequality $\alpha_{nk}>\beta_{nk}$. Therefore, the wavefunction-continuity conditions at the point of barrier-well junction, such as $x=0$, can be written as:
@@
According to the results of the KP model, we find that the PBA of quantum state is dramatically consistent with the classical physical intuition: a free particle has an initial kinetic energy $T_{0}$; When it passes through the barrier region $V(x)>0$, part of the kinetic energy is converted into potential energy, and its velocity decreases due to the smaller kinetic energy $T(x)=T_{0}-V(x)<T_{0}$, causing it to have a higher probability of being detected in the higher barrier region. When the initial kinetic energy is much higher than the potential barrier $T_{0}>>V(x)$, the particle can move freely throughout space $T(x)\approx T_{0}$. This finding seems to indicate that high-energy PBA electrons tend to classical behavior according to Bohr’s correspondence principle[31].
@@ Methods
Figure 5: The Kohn-Sham effective potential and electron density of Na+-hP4 in (a) (110) plane, and (b) $\bar{1}1$] direction. The grey line denotes the reference potential of Na-hP4.
To isolate the interatomic potential field from the screening effect of valence electrons, we calculated the KSDFT-derived effective potential for a hypothetical system of Na+-hP4 by removing the valence electrons out of Na-hP4. As shown in Figure 5(a) and (b), the maximum of this “ion-only” potential remains located at the interstitial site. This is expected, as the potential converges toward the superposition of screened ionic Coulomb potentials in the interatomic region. Notably, the absence of near-free electron screening leads to a higher interatomic potential barrier in Na+-hP4 than in Na-hP4. Crucially, this independent calculation confirms our central thesis: interstitial anionic electrons (IAEs) are near-free electrons with high density accumulation around the potential maximum, not bound states confined to potential wells.
@@
Generalized doulbe-barrier Kronig-Penney model
Figure 6: (a) The effective potential and electron density calculated by the all-electron potential for Diamond; (b) The generalized KP model with two different barriers (or wells) in a unit cell. “A”/“C” and “B”/“D” denote the interatomic regions and atomic regions, respectively. Especially, “C” also indicates the covalent bonding region.
Figure 4(c) of the maintext shows that the effective potential in diamond reaches a maximum at the carbon sites. This feature originates from the pseudopotential representing the core-electron interaction, an inherent characteristic of the frozen-core approximation within the PAW method. To clarify the physical picture, we performed all-electron (AE) calculations, the results of which are presented in Figure 6(a). These calculations confirm the highly localized core-electron density at the carbon nuclei and the corresponding deep effective potential wells centered on these atomic sites.
To give a physical picture of the covalent bonding [see Figure 4(c) in our maintext], we further propose a more generalizable extension of the Kronig-Penney (KP) model, which can be implemented as a double-barrier (or double-well) in a unit cell for the diamond structure, as illustrated in Figure 6(b). Here, we shift the maximum potential to 0 eV for convenience.
For such a square barrier-well potential, the components of the wavefunction adopt one of two different forms depending on the electron energy: (i) an oscillatory, planewave-like solution when the energy exceeds the potential barrier, or (ii) an exponentially decaying solution when the energy is below the potential barrier. For the square barrier-well potential with the configuration of $V_{B}\leq V_{D}<V_{C}<V_{A}$ as shown in Figure 6(b), the energy levels ($\epsilon_{nk}$) can be classified into three categories: (1) bound states ($V_{B}<\epsilon_{nk}<V_{C}$) for localized electrons (core states); (2) semibound states ($V_{C}<\epsilon_{nk}<V_{A}$) for covalent bonds; and (3) unbound states ($\epsilon_{nk}>V_{A}$) for near-free electrons, metallic bonds, or interstitial anionic electrons.
For comparison with the covalent bonds in the diamond system, our analysis focuses on the semibound state in Figure 6(b) with energy in between the two barriers ($V_{C}<\epsilon_{nk}<V_{A}$ ), where the wavefunction components in different regions can be written as:
$\psi_{nk}^{\text{A}}(x)=A_{nk}e^{\alpha_{nk}x}+A^{\prime}_{nk}e^{-\alpha_{nk}x},\penalty 10000\ x\in\text{A};$
$\psi_{nk}^{\text{B}}(x)=B_{nk}e^{i\beta_{nk}x}+B^{\prime}_{nk}e^{-i\beta_{nk}x},\penalty 10000\ x\in\text{B};$
$\psi_{nk}^{\text{C}}(x)=C_{nk}e^{i\gamma_{nk}x}+C^{\prime}_{nk}e^{-i\gamma_{nk}x},\penalty 10000\ x\in\text{C};$
$\psi_{nk}^{\text{D}}(x)=D_{nk}e^{i\eta_{nk}x}+D^{\prime}_{nk}e^{-i\eta_{nk}x},\penalty 10000\ x\in\text{D}.$ (11)
We emphasize that in region A, the wavefunction must decay exponentially, as indicated by the Schrödinger equation for a classically forbidden region. This behavior is fundamentally different from the planewave-like character of the wavefunctions in regions B, C, and D. With the eight numerical conditions (six continuity conditions at the points of $x=\{x_{1},x_{2},x_{3}\}$; two periodic boundary conditions at $x=0$ and $x=x_{4}$), the eight coefficients in Eq. (11) can be strictly solved. The potential-barrier affinity effect can occur in region C with planewave-like wavefunctions, as demonstrated in the maintext of our manuscript. We note that systematically adjusting the relative height and spatial separation of the potential barriers in this generalized model may yield deeper insights into the understanding of the formation mechanisms of different bonding types. For the AE KSDFT calculations, we performed the calculations using the full-potential linearised augmented-planewave method[37] by the ELK-v10.6.2 code[38]. $18\times 18\times 18$ $k$-grid meshes are used for the ground-state calculations of the Diamond crystal. The potential and electron density are represented in the dense real-space grids of $120\times 120\times 120$.